Four less than half a number is equal to 37 less than twice the number. How do you Translate the sentence into an equation and Solve the equation.

Answer by stanbon(75887)

Four less than half a number is equal to 37 less than twice the number. How do you Translate the sentence into an equation and Solve the equation.
-------
(1/2)x - 4 = 2x-37
------
(3/2)x = 33
----
(1/2)x = 11
x = 22
-------------
Cheers,
Stan H.
---------------